An aquarium filled with water has flat glass sides (thickness D = 1.25 cm). A beam of light from outside the aquarium strikes the glass at angle \Theta = 53.1

What's the angle (relative to the perpendicular to the surface) of this light ray when it enters

(a) the glass (nglass = 1.50)?
(b) the water (nwater = 1.33)?
(c) Find the speed of light in glass
(d) Find the time it takes the light ray to travel through the glass plate
Answer #1

D = 1.25 cm, let theta be x

x = 53.1 ,
we know n1 sin i = n2 sin r
here n1 = nair = 1, and i = x = 53.1
n2 = nglass = 1.5,
so 1 * sin 53.1 = 1.5 * sin r
=> sin r = 0.533
=> r = 32.2

a) angle when it enters glass = 32.2

now when going from glass to water we again have

nglass sin i = nwater sin r
1.5 * sin 32.2 = 1.33 * sin r
=> sin r = 0.6
=> r = 36.87

b) angle when it enter water = 36.87

we know speed of light be vmedium

vglass nglass = vair nair
vglass = 1 * 3 *108 / 1.5 = 2 * 108 m/s

c) speed of light in glass = 2 * 108 m/s

time it will take = D/vglass = 1.25 * 10-2 m / 2 * 108 m/s = 6.25 * 10-11 s

d) time it takes to travel through glass = 6.25 * 10-11 s
Answer #2

Answer question